OpenStudy (anonymous):

14^(x+1)=36 (round to nearest ten thousandth)

OpenStudy (campbell_st):

take the log of both sides either base e or base 10 \[\ln(14^{x +1}) = \ln(36)\] so apply the log law for powers... and you get \[(x + 1)\ln(14) = \ln(36)\] now solve for x

OpenStudy (anonymous):

x is 0.36

OpenStudy (campbell_st):

looks good to me... well done